NDOS Commissioning Meeting 2011/08/18

Attendance: Jon, Susan, Jan, Martin, Rafael, Leon, Minerba, Mat, Athans, Gary, Jarek, Mark Messier, Peter S., Andrew


  1. Beam & Detector Update
  2. DataCheck Update
  3. DAQ Update
  4. DCS Update
  5. AOB

Beam & Detector Update

  • Of the 20 APDs that had cooling enabled last week, 4 had already been removed from the detector but not from the golden channel list. Of these 16, 6 have failed, been removed by Mat and shipped off to CalTech.
  • Another request to cool more APDs, those that have been shimmed, in order to test if the failures are occurring
  • Jon spoke with Chuck Andrews to see if he is available to help us debug the possible grounding problem on the NDOS. Chuck was busy this week, but should now be available to help out. Jon will follow up.
  • Leon:
    • got first version of software today that will allow for overnight tests of the APDs; will begin testing today/tonight (can do 14 at a time)
    • Peter: would it be helpful to have a DAQ system at CalTech? Leon: probably, yes.
    • Jon: any further progress in recovering other APD? Yes, but it's slow going.

DataCheck Update

  • Still not clear why DataCheck metrics do not show a problem with the DCM timing seen this past week.
  • Cosmic ray track reconstruction is possibly the best way to look for these kinds of timing problems.
  • Some progress in getting the nearline processing going.

DCS Update

  • Martin:
    • getting GUI to interface to dB in order to allow one to control cooling from the GUI
    • power script testing tomorrow? Jon, Martin and Peter discussed schedules and Martin will do tests at 9 am tomorrow.

DAQ Update

  • Peter: in addition to gross mis-synchronization found last week, another mis-synchronization was found where dcms are hundreds of micro-seconds off. This has been tracked down to a mis-ordering of commands. The order has been reversed after which the clocks are synchronized. RC will need to wait ~2.2s before issuing a BeginRun to GlobalTrigger.